Bag A contains 25 tomatoes, 4 of which are rotten. Bag B contains 15 tomatoes, 6 of which are rotten. Jamie randomly selects a tomato from bag A followed by another random selection from bag B. What is the probability that exactly one tomato is rotten? What is the probability that at least one tomato is not rotten?

Answer by greenestamps(13373)



Given:
P(A rotten) = 4/25 --> P(A not rotten) = 21/25
P(B rotten) = 6/15 --> P(B not rotten) = 9/15

(a) P(exactly one is rotten) = P((A rotten) AND (B not rotten) OR (A not rotten) AND (B rotten))

(b) P(at least one not rotten) = 1-P((A rotten) AND (B rotten))

Plug in the given probabilities, using multiplication for AND and addition for OR.